OpenStudy (anonymous):

11. Solve the inequality. Show your work. |r + 3| ≥ 7

OpenStudy (dtan5457):

Absolute value means it can go 2 ways \[r+3\ge7\] \[r+3\le-7\]

OpenStudy (dtan5457):

Can you solve for r?

OpenStudy (anonymous):

I am not good at algebra at all so I dont know.

OpenStudy (dtan5457):

Isolate r. \[r+3\ge7\] Subtract 3 \[r \ge4\]

OpenStudy (dtan5457):

Can you try for the 2nd one..

OpenStudy (anonymous):

subtract 3 from what 7-3

OpenStudy (dtan5457):

-7-3=?

OpenStudy (anonymous):

-10

OpenStudy (dtan5457):

\[r \le-10\]

OpenStudy (anonymous):

so -7-3=-10 so that answer would be r<-10

OpenStudy (dtan5457):

yes

OpenStudy (anonymous):

Is that all I have to do

OpenStudy (dtan5457):

yep

OpenStudy (anonymous):

What percent of 60 is 12? Show your work. Alright is think I got this but not sure. I did 60%x12=7.2 is that right